Find the x intercepts of y = x^2 - 4x
Note x^2 should be read as x squared

Found 2 solutions by richwmiller, ewatrrr:
Answer by richwmiller(17219)

0=x*(x-4)
x=0 x=4

Answer by ewatrrr(24785)

 
Hi,
Find the x intercepts of
y = x^2 - 4x
y = x(x - 4) |Note: y = 0 when curve crosses x -axis
x(x-4) = 0 x = 0 0r x = 4
Curve crosses the x-axis at (0,0) and (4,0)